Suppose that a chemistry student can obtain two acid solutions from the stock room. The first solution is 20% acid and the second solution is 45% acid. (The percentages are by volume.) How many milliliters of each solution should the student mix together to obtain 100 milliliters of 30% acid solution?
Answer by mananth(16949)

percent ---------------- quantity
Acid solution I 20.00% ---------------- x ml
Acid solution II 45.00% ------ 100 - x ml
Mixture 30.00% ---------------- 100
12
20.00%x+45.00%(100-x)= 30.00% *100
20x+45 (100-x) =3000
20x+4500-45x=3000
20x-45x=3000-4500
-25x= -1500
/ -25
x=60 ml 20.00% Acid solution I
40 ml 45.00% Acid solution II

m.ananth@hotmail.ca